OpenStudy (anonymous):

a positive integer m divides 748x+714y for all values of x and y .how many values of m can we have ?

OpenStudy (anonymous):

linear combination of 748 and 714

OpenStudy (anonymous):

:O , i didnt understand

OpenStudy (anonymous):

4 values of m

OpenStudy (anonymous):

can u explain how you got this?

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

i dont know but the options are :- 1)2 2)3 3)4 4) infinitely many

OpenStudy (anonymous):

748x+714y=748(1)+(-1)714

OpenStudy (anonymous):

m can be 1,2,17,34

OpenStudy (anonymous):

firstly note that the greatest common divisor of 748 and 714 is 34 because: \[748=2^2*11*17\]\[714=2*3*7*17\]now we know that m divides\[34(22x+21y)\]for all values of x and y specially for x=1 and y=-1 which will gives the number as\[34(22-21)=34\]so m must be a divisor of 34 and since we know 34=2*17 so\[m=1,2,17,34\]

OpenStudy (anonymous):

we let x=1 and y=-1 because we want to make the 22x+21y equal to 1

OpenStudy (anonymous):

thanks ! it really helps :)

OpenStudy (anonymous):

yw :)